OpenStudy (anonymous):

1) Select the BEST classification for π. A) irrational B) rational C) imaginary D) complex

OpenStudy (mathstudent55):

First, find the definitions of those terms. Then read them and understand them. Finally, pick the correct answer. If you still need help, then ask.

zepdrix (zepdrix):

Hmm I'm trying to think of how to give you a hint without just giving away the answer >.< Well here are some examples of numbers: Rational can be written as a ratio of whole numbers:\[\large\rm \frac{2}{3},\qquad\qquad \frac{7}{11},\qquad\qquad \frac{15}{22}\] Irrational numbers have no fractional representation in the whole numbers:\[\large\rm 4\pi,\qquad\qquad \sqrt{3},\qquad\qquad \sqrt{7}\] Imaginary is a number undefined in the real numbers:\[\large\rm \sqrt{-1},\qquad\qquad \sqrt{-4},\qquad\qquad \sqrt{-12}\] Complex numbers have both a real and imaginary part:\[\large\rm \pi+\sqrt{-2},\qquad\qquad 7-\sqrt{-5},\qquad\qquad 11+\sqrt{-11}\]

zepdrix (zepdrix):

So I gave some examples of each type of number. Any ideas which group pi would belong to? :)

OpenStudy (anonymous):

imaginary

zepdrix (zepdrix):

Hmm

OpenStudy (anonymous):

i have another problem i cant answer Seats in an auditorium are numbered 1-500, and the principal is handing out snacks in an assembly. Every student whose seat is a multiple of 4 gets chocolate. Every student in a seat that is a multiple of 7 gets jelly beans, and every student in a seat that is a multiple of 10 gets chips. The student in which seat number is the first to get all 3 treats? A) 40 B) 70 C) 100 D) 140

OpenStudy (mathstudent55):

Here are some important definitions: Natural numbers: the counting numbers: 1, 2, 3, 4, ... Whole numbers: the counting numbers and 0: 0, 1, 2, 3, 4, ... Integers: the whole numbers and the negatives of the whole numbers: ... , -3, -2, -1, 0, 1, 2, 3, ... Rational numbers: any number that can be written as a fraction of integers. Rational numbers either terminate: 2, 2,.55, 4.301, or they don't terminate, but they repeat: 3.555..., 4.636363... Irrational numbers: numbers that cannot be written as a fraction of integers. Irrational numbers do not terminate and do not repeat. All numbers described up to here form the set of real numbers. Imaginary: a number that has as a factor the square root of -1, denoted as i. Complex number: a number that is a sum of a real part and an imaginary part.